OpenStudy (anonymous):

Solve 8 + √2x less than or equal to 5?

zepdrix (zepdrix):

Hey Alexis :) Is the x under the root like this? \(\large\rm 8+\sqrt{2x}\le 5\)

OpenStudy (anonymous):

Yes :)

zepdrix (zepdrix):

We have a bunch of `operations` being applied to x. We need to undo all of them in order to isolate the x. We have some addition, some multiplication, some square root. We'll undo all of that by starting with the most basic. We have an 8 being added, how do we undo that? :O Reverse of addition?

OpenStudy (anonymous):

We would subtract 8

OpenStudy (anonymous):

From both sides

zepdrix (zepdrix):

Mmm k good. That's a good start.\[\large\rm 8+\sqrt{2x}-8\le 5-8\]\[\large\rm \sqrt{2x}\le -3\]Hmm the multiplication is `inside` of the root, so we can't get to that operation quite yet. How do we deal with the square root? What is the opposite of square rooting? :O

OpenStudy (anonymous):

We would square the number by itself to isolate it out?

zepdrix (zepdrix):

Good, squaring is the inverse of square rooting. Think back to what you did to undo addition. You applied subtraction `to both sides`. We'll simply do that with this step as well. We'll `square both sides`.

zepdrix (zepdrix):

\[\large\rm \left(\sqrt{2x}\right)^2\le \left(-3\right)^2\]So on the left side, the square root and the square `undo` one another, In the same way that +8 and -8 undo one another. On the right side, squaring -3 gives us (-3)(-3)=9.\[\large\rm 2x\le 9\]

zepdrix (zepdrix):

Last step?? :)

OpenStudy (anonymous):

That is where I am stuck...

zepdrix (zepdrix):

Even though you don't see a symbol, the 2 is `multiplying` the x. You need an operation that will undo multiplication.

OpenStudy (anonymous):

Divide both sides by 2?

zepdrix (zepdrix):

Good! :)\[\large\rm \frac{\cancel2x}{\cancel2}\le \frac{9}{2}\]Giving us our final result of\[\large\rm x\le \frac{9}{2}\]

zepdrix (zepdrix):

Yayyy team \c:/

OpenStudy (anonymous):

Omg! thank you so much!!!! I really appreciate your effort and time!